# Manolis Mylonas CERTH-ITI Thessaloniki, Greece

### v1

@@ 5.2 Implementation details
Similarly to [9], videos are represented using one frame per second, and frame-level embeddings of the visual content (of size $D=512$) are obtained using a fine-tuned CLIP model on the data of VideoXum, that has been released by the authors of [9] [^2] Publicly-available at: https://videoxum.github.io/. The same model is used to extract sentence-level embeddings (of the same size $D=512$) from the generated scripts. With respect to the SD-VSum network architecture, the number of attention heads is set equal to $8$, and the Transformer-based frame scoring mechanism consists of a Transformer network [34], a linear layer with $512$ neurons, and a sigmoid function that computes the frame-level importance scores. The learning rate, dropout rate and L2 regularization factor are equal to $5\cdot 10^{-5}$, $0.5$ and $10^{-4}$, respectively. For the network’s weights initialization, we use the Xavier uniform initialization approach with gain = $\sqrt{2}$ and biases = $0.1$. Training is performed in a batch mode with a batch size equal to $4$ using the Adam optimizer, and stops after $50$ epochs. All experiments were carried out on a PC with an NVIDIA RTX 3090 GPU. To allow the reproduction of our results, the used data and the PyTorch implementation of SD-VSum will be made publicly-available upon paper’s acceptance.
5.3 Experimental comparisons and ablations
@@
Table 2: Script-driven video summarization: performance comparison (F-Score (%)) with the CLIP-It method and several variants of the proposed SD-VSum network architecture.
@@
CLIP-It [6] single vector 4 Yes 22.8
@@
The proposed SD-VSum network architecture was compared with CLIP-It [6] that exhibits state-of-the-art performance on query-driven and multimodal (generic) video summarization. The results in the upper part of Table 2 show that SD-VSum performs clearly better than CLIP-It on the extended S-VideoXum dataset, achieving an F-Score that is almost $2\%$ higher than the F-Score of CLIP-It. Since SD-VSum and CLIP-It are similar in terms of network architecture (as discussed in Section 4.1), we tried to spot what causes this noticeable difference in the script-driven summarization performance of these architectures. For this, we conducted an ablation study that involved variants of SD-VSum that resemble the network architecture of CLIP-It, as well as a variant that focuses on the use of a scaling process when computing the attention weights. In particular, we considered the following variants:
@@
  • Variant #2 extends the previous variant by containing the same number of attention heads with CLIP-It.
  • Variant #3 extends the previous variant by performing a scaling process when computing the attention weights (by dividing with $\sqrt{D}$, where $D$ is the size of the used embeddings), simulating further the language-guided attention mechanism of CLIP-It.
@@
The findings of our ablation study are reported in the lower part of Table 2. The performance of Variant #1 highlights the impact of producing and using a significantly more condensed representation for the user script, which leads to a noticeable drop in the script-driven video summarization performance (almost $2\%$ lower than SD-VSum). The performance of Variant #2 indicates the slight effect of the scaling process when using a single vector representation for the user script ($0.1\%$ lower than the performance of Variant #1), an effect that is more pronounced when the script is represented using multiple embeddings (as denoted by the performance of Variant #4 which is $0.6\%$ lower than that of SD-VSum). Finally, the reduction of the number of employed attention heads leads to further performance reduction ($0.5\%$ lower than the performance of Variant #2). In general, the observed performance for the CLIP-It method and the variants of SD-VSum that utilize the same approach for representing the user script, documents the contribution of using a more coarse-grained sentence-level representation of the user script and points out the effectiveness of the integrated cross-modal attention mechanism in the SD-VSum network architecture.
Table 3: Performance comparison between the proposed SD-VSum network architecture for script-driven video summarization and literature networks for visual-based and multimodal video summarization, in terms of F-Score (%) and Kendall’s $\tau$ and Spearman’s $\rho$ rank correlation coefficients.
@@
Script-driven
@@
SD-VSum 24.8 N/A N/A
Multimodal summarization SD-VSum 23.6 0.188 0.247
VTSUM-BLIP [9] 23.5 0.196 0.258
CLIP-It [6] 22.8 0.154 0.203
Visual-based summarization VSUM-BLIP [9] 23.1 0.185 0.246
PGL-SUM [30] 22.0 0.153 0.203
Our next experiments focused on assessing the capacity of generic video summarization methods to tackle the needs of the script-based video summarization task. For this, we evaluated the performance of SD-VSum and CLIP-It after using as input a natural language description of the full-length video (instead of the desired video summary). Moreover, we ran experiments with a state-of-the-art method for visual-based video summarization (PGL-SUM [30]) that estimates the frames’ importance using multi-head attention mechanisms that model the frames’ dependence at different levels of granularity. As a note, the computation of the Kendall’s $\tau$ and Spearman’s $\rho$ rank correlation coefficients in the case of script-driven video summarization is not possible, since the output of the network architecture is compared with a binary ground-truth summary. Our comparisons included also the visual-based and multimodal network architectures that were evaluated by the authors of the VideoXum dataset, after taking into account the scores reported in [9]. The outcomes of this investigation, presented in Table 3, show that the visual-based video summarization approaches are the least effective ones; thus, using a description of the full-length video as auxiliary input information, seems to be beneficial for the summarization process. Concerning the multimodal approaches, the performance of SD-VSum is clearly better than the one of CLIP-It, and highly competitive to the one of VTSUM-BLIP. Nevertheless, it is lower than the exhibited performance when SD-VSum is used for script-driven video summarization. These findings indicate the importance of script-driven video summarization technologies, and their competency to produce video summaries that are better tailored to the needs of different users.
@@
Table 4: Comparison in terms of computational complexity.
@@
CLIP-It [6] 26.17 45.98
5.4 Qualitative analysis
@@
Our qualitative analysis aimed to investigate the extend to which the output of script-driven video summarization is more tailored to the needs of different users, compared to the generic video summaries that are produced by multimodal approaches. For this, we compared the overlap (expressed in F-Score (%)) between the multiple available ground-truth summaries for $50$ videos of the extended S-VideoXum dataset, and the produced summaries when the proposed SD-VSum is used for script-driven and multimodal video summarization, respectively. The pool of videos was defined by choosing the ones that showed the major diversity in the computed overlap when comparing the multimodal-based video summary with each different ground-truth summary, and thus were the most challenging ones for a multimodal video summarization approach. The obtained scores, that are illustrated in the form of heatmap values in Fig. 4, demonstrate that the output of multimodal video summarization (see the heatmap on the right side) shows the same weak overlap with the summaries of the different human annotators, and thus is poorly adapted to different needs about the content of the summary. On the contrary, the output of script-driven video summarization is clearly more aligned to the preferences of each different annotator, for most of the selected videos. These findings justify once again the motivation behind the development and training of network architectures for script-driven video summarization, and highlight the capacity of the proposed SD-VSum network architecture to produce video summaries that are well-tailored to each user’s needs about their content.
@@
Figure 4: The overlap (expressed in F-Score (%)) between the multiple available ground-truth summaries for $50$ videos of the extended S-VideoXum dataset, and the produced video summaries by SD-VSum when it is used for script-driven (left side) and multimodal video summarization (right side).
6 Conclusions
@@
In this paper, we introduced a new task that aims to create a video summary of the full-length video by choosing the parts showing the highest relevance with a user-provided script that outlines the content of the desired summary. To support our research, we extended the VideoXum large-scale dataset for video summarization by generating natural language descriptions of the different ground-truth summaries that are available per video. Using the extended S-VideoXum dataset, we trained and evaluated a new network architecture for script-driven video summarization (SD-VSum), that integrates a cross-modal attention mechanism to align and fuse information from the different data modalities. Our quantitative evaluations highlighted the advanced performance of SD-VSum compared to state-of-the-art methods for query-driven and generic (unimodal and multimodal) summarization from the literature, and pointed out the impact of using coarse-grained representations for the input text. Our qualitative analysis demonstrated the capacity of SD-VSum to incorporate knowledge about the user’s needs and produce well-customized video summaries.

### v2

@@ 5.2. Implementation details
Similarly to (Lin et al., 2024), videos are represented using one frame per second, and frame-level embeddings of the visual content (of size $D=512$) are obtained using a fine-tuned CLIP model on the data of VideoXum, that has been released by the authors of (Lin et al., 2024) [^3] Publicly-available at: https://videoxum.github.io/. The same model is used to extract sentence-level embeddings (of the same size $D=512$) from the generated scripts. Following the paradigm of e.g., CLIP-It (Narasimhan et al., 2021), we keep both the visual and the text encoder frozen. With respect to the SD-VSum network architecture, the number of attention heads is set equal to $8$, and the Transformer-based frame scoring mechanism consists of a Transformer network (Vaswani et al., 2017), a linear layer with $512$ neurons, and a sigmoid function that computes the frame-level importance scores. The learning rate, dropout rate and L2 regularization factor are set equal to $5\cdot 10^{-5}$, $0.5$ and $10^{-4}$, respectively. For the network’s weights initialization, we use the Xavier uniform initialization approach with gain = $\sqrt{2}$ and biases = $0.1$. Training is performed in a batch mode with a batch size equal to $4$ using the Adam optimizer, and stops after $50$ epochs. All experiments were carried out on a PC with an NVIDIA RTX 3090 GPU. The PyTorch implementation of SD-VSum is publicly-available at: https://github.com/IDT-ITI/SD-VSum.
5.3. Experimental comparisons and ablations
The proposed SD-VSum network architecture was compared with CLIP-It (Narasimhan et al., 2021), that exhibits SOTA performance on query-driven and multimodal (generic) video summarization. The results in the upper part of Table 2 show that SD-VSum performs clearly better than CLIP-It on the extended S-VideoXum dataset, achieving an F-Score that is almost $2\%$ higher than the F-Score of CLIP-It. A similar performance gap between SD-VSum and CLIP-It was observed on the small-scale S-NewsVSum dataset; the computed F-Scores for SD-VSum and CLIP-It were $22.4$ and $20.0$, respectively.
Given the diversity of the generated scripts, we analyzed the performance of SD-VSum across varying script lengths using four subsets of S-VideoXum: S1: $10\%$ of scripts with fewest words (mean #words: $66.8$); S2: $10\%$ of scripts with fewest sentences (mean #sentences: $3.8$); S3: $10\%$ of scripts with most words (mean #words: $157.2$); S4: $10\%$ of scripts with most sentences (mean #sentences: $8.9$). We observed very small performance fluctuations across subsets S1 to S4 (F-Score = $24.5$, $24.5$, $24.7$, $25.0$, respectively); SD-VSum performs slightly better for scripts made of many sentences.
@@ 5.3. Experimental comparisons and ablations
Table 2. Script-driven video summarization: performance comparison (F-Score (%)) with the CLIP-It method and several variants of the proposed SD-VSum network architecture.
@@
CLIP-It (Narasimhan et al., 2021) single vector 4 Yes 22.8
@@
Following, since SD-VSum and CLIP-It are similar in terms of network architecture, we tried to spot what causes the observed difference in their performance. For this, we ran an ablation study including the following variants of SD-VSum that resemble the network architecture of CLIP-It, as well as a variant that performs a scaling process when computing the attention weights:
@@
  • Variant #2 extends Variant #1 by containing the same number of attention heads with CLIP-It.
  • Variant #3 extends Variant #2 by performing a scaling process when computing the attention weights (by dividing with $\sqrt{D}$, where $D$ is the size of the used embeddings), simulating further the language-guided attention mechanism of CLIP-It.
@@
Focusing on the lower part of Table 2, the performance of Variant #1 shows that the use of a significantly more condensed representation for the user script leads to a noticeable drop in the summarization performance ($2\%$ lower than SD-VSum). The performance of Variant #2 indicates the slight effect of the scaling process when using a single vector representation for the user script ($0.1\%$ lower than that of Variant #1), an effect that is more pronounced when the script is represented using multiple embeddings, as denoted by the performance of Variant #4 which is $0.6\%$ lower than that of SD-VSum. Finally, the use of fewer attention heads leads to further performance reduction ($0.5\%$ lower than the performance of Variant #2). These findings document that the use of multiple sentence-level embeddings for text representation (a change compared to CLIP-It) by the integrated cross-modal attention mechanism is critical for addressing the new problem of script-driven summarization, and demonstrate the competency of SD-VSum to deal with scripts describing multiple unrelated actions.
Subsequently, we assessed the capacity of generic video summarization methods to tackle the script-driven video summarization task. For this, we evaluated the performance of SD-VSum and CLIP-It using as input a textual description of the full-length video. Moreover, we ran experiments with a SOTA method for visual-based video summarization (PGL-SUM (Apostolidis et al., 2021b)) which estimates frames’ importance after modeling their dependence at different levels of granularity using multi-head attention mechanisms. As a note, the computation of the Kendall’s $\tau$ and Spearman’s $\rho$ rank correlation coefficients for script-driven video summarization is not possible, since the network’s output is compared with a binary ground-truth summary. Our comparisons included also the visual-based and multimodal network architectures evaluated by the authors of the VideoXum dataset, using the scores reported in (Lin et al., 2024). The outcomes of this investigation (see Table 3), show that the visual-based approaches are the least effective ones; thus, using a description of the full-length video as auxiliary input information seems to be beneficial for the summarization process. Concerning the multimodal approaches, the performance of SD-VSum is clearly better than that of CLIP-It, and highly competitive to the one of VTSUM-BLIP. Nevertheless, it is lower than the observed performance when SD-VSum is used for script-driven video summarization. These findings indicate the importance of script-driven video summarization technologies, and their capacity to produce video summaries that are better tailored to the needs of different users.
Table 3. Performance comparison between the proposed SD-VSum network architecture for script-driven video summarization and literature methods for visual-based and multimodal video summarization, in terms of F-Score (%) and Kendall’s $\tau$ and Spearman’s $\rho$ rank correlation coefficients.
@@
Script-driven SD-VSum 24.8 N/A N/A
Multimodal
@@
SD-VSum 23.6 0.188 0.247
VTSUM-BLIP (Lin et al., 2024) 23.5 0.196 0.258
CLIP-It (Narasimhan et al., 2021) 22.8 0.154 0.203
Visual-based
summarization
VSUM-BLIP (Lin et al., 2024) 23.1 0.185 0.246
PGL-SUM (Apostolidis et al., 2021b) 22.0 0.153 0.203
@@
Table 4. Comparison in terms of computational complexity.
@@
CLIP-It (Narasimhan et al., 2021) 26.17 45.98
5.4. Qualitative analysis
@@
Our qualitative analysis aimed to investigate the extent to which the output of script-driven video summarization is more tailored to the needs of different users, compared to summaries produced by multimodal approaches. For this, we compared the overlap (expressed in F-Score (%)) between the multiple available ground-truth summaries for $50$ videos of the S-VideoXum dataset, and the produced summaries when SD-VSum is used for script-driven and multimodal video summarization, respectively. The pool of videos was defined by choosing the ones showing the major diversity in the computed overlap when comparing the multimodal-based video summary with each different ground-truth summary, and thus were the most challenging ones for a multimodal video summarization approach. The obtained scores, that are illustrated in the form of heatmap values in Fig. 4, demonstrate that the output of multimodal video summarization (see the heatmap on the right side) shows the same weak overlap with the summaries of the different human annotators, and thus is poorly adapted to different needs about the content of the summary. On the contrary, the output of script-driven video summarization is clearly more aligned to the preferences of each different annotator, for most of the selected videos. These findings justify once again the motivation behind the development and training of network architectures for script-driven video summarization, and highlight the competency of the proposed SD-VSum network architecture to produce video summaries that are well-tailored to each user’s needs about their content.
@@
Figure 4. The overlap (expressed in F-Score (%)) between the multiple available ground-truth summaries for $50$ videos of the extended S-VideoXum dataset, and the produced video summaries by SD-VSum when it is used for script-driven (left side) and multimodal video summarization (right side).
6. Conclusions
@@
In this paper, we introduced a new task that aims to create a video summary of the full-length video by choosing the parts showing the highest relevance with a user-provided script that outlines the content of the desired summary. To support our research, we extended the VideoXum large-scale dataset for video summarization by generating natural language descriptions of the different ground-truth summaries that are available per video. Using the extended S-VideoXum dataset, we trained and evaluated a new network architecture for script-driven video summarization (SD-VSum), that integrates a cross-modal attention mechanism to align and fuse information from the different data modalities. Our quantitative evaluations highlighted the advanced performance of SD-VSum compared to SOTA methods for query-driven and generic (unimodal and multimodal) video summarization from the literature, and pointed out the impact of using coarse-grained representations for the input text. Our qualitative analysis demonstrated the capacity of SD-VSum to incorporate knowledge about the user’s needs and produce well-customized video summaries.
Acknowledgements.
This work was supported by project MediaPot ($TAE\Delta K$-06196), carried out within the framework of the National Recovery and Resilience Plan Greece 2.0, funded by the European Union – NextGenerationEU, under the call RESEARCH-CREATE-INNOVATE.